OpenStudy (deathlynine):

Pam laid three times as many tiles on Saturday as she laid on Sunday. If Pam laid 80 tiles over the weekend altogether, how many tiles did she lay on Saturday?

OpenStudy (prathamesh_m):

Let the number of tiles laid on Sunday be x. Then the number of tiles laid on Saturday will be 3x. The sum of those two would be the total number of tiles, i.e., 80.

OpenStudy (prathamesh_m):

So once you find x, 3x would be the number of tiles laid on Saturday.

OpenStudy (deathlynine):

So the equation would be x + 3x = 80?

OpenStudy (prathamesh_m):

Yes that's right.

OpenStudy (deathlynine):

Can you check my answer? x + 3x = 80 4x = 80 4x/4 = 80/4 20 <----- I got this, is this correct?

OpenStudy (prathamesh_m):

Yes that's the value of x.

OpenStudy (deathlynine):

ok thanks for helping me. :)

OpenStudy (prathamesh_m):

np